i have a rectangular metal frame. The sides are 6 feet tall by 12 foot long horizontal pieces. what is the diagonal measurement of this rectangle?
Answer by checkley77(12569)

6^2+12^2=X^2
36+144=X^2
180=X^2
X=SQRT180
X=13.4164 ANS. FOR THE DIAGONAL.